OpenStudy (anonymous):

PLEASE HELP!! write an equation in slope-intercept form for the line that passes through (0,6) and is parallel to the line given by y=2x+3 .

OpenStudy (ashley1nonly):

first sorry for my spelling

OpenStudy (anonymous):

general linear equation y=mx+c m=gradient c=y-intercept perpendicular tells us that we have to use this formula m1 x m2 = -1 m1=gradient of equation given m1=2 m2=gradient of unknown equation we want to find 2 x m2 = -1 m2=-1/2 now substitute (4,6) and gradient m=-1/2 into the general equation y=mx+c to find the c value and then re-write the whole equation.

OpenStudy (ashley1nonly):

are we looking for a prendicular line or parrallel

OpenStudy (anonymous):

perpendicular line, and this helps us figure out the unknown equation perpendicular to the one given.

OpenStudy (ashley1nonly):

if it is parrarllel then the slope is the same but prependicular you would use the negative recipcoral

OpenStudy (anonymous):

yes that is correct.

OpenStudy (ashley1nonly):

in the equation he says that is is parralel to y=2x+3 so why did you solve for the prendicular form you could have used y-y1=m(x-xx1) you already have the slope which is 2

OpenStudy (ashley1nonly):

then wouldnt you plug in the points you have and then slove

OpenStudy (ashley1nonly):

0,6 y-6=2(x-0) y-6=2x y=2x+6

OpenStudy (ashley1nonly):

used my graphing calc and they were both parralel in a diganol way

OpenStudy (ashley1nonly):

medal if right

OpenStudy (ashley1nonly):

did yiu have any questions

OpenStudy (anonymous):

no thanks

OpenStudy (ashley1nonly):

okay